OpenStudy (anonymous):

Write the expression as the sine, cosine, or tangent of an angle. cos 107° cos 42° + sin 107° sin 42°

OpenStudy (anonymous):

These are the answers sin 65° cos 65° cos 149° sin 149°

OpenStudy (anonymous):

No this is not about triangles I'm afraid.

OpenStudy (anonymous):

just use the identities like \[\huge \cos(A-B)=\cos A \cos B + \sin A \sin B\]

OpenStudy (anonymous):

So cos(107-42)?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Yay which equals 65! Thanks!

OpenStudy (anonymous):

welcome!